OpenStudy (anonymous):

Find the functional values requested in the problem. h(x)=2x-1 (a) h(0)= (b) h(-6)=

OpenStudy (anonymous):

(a) h(0) = 2*0 -1 = 0-1 = -1 (b) h(-6) = 2*-6 -1 = -12 -1 = -13